# DLTPose: 6DoF Pose Estimation From Accurate Dense Surface Point Estimates

### v1

@@ Abstract
We propose DLTPose, a novel method for 6DoF object pose estimation from RGBD images that combines the accuracy of sparse keypoint methods with the robustness of dense pixel-wise predictions. DLTPose predicts per-pixel radial distances to a set of minimally four keypoints, which are then fed into our novel Direct Linear Transform (DLT) formulation to produce accurate 3D object frame surface estimates, leading to better 6DoF pose estimation. Additionally, we introduce a novel symmetry-aware keypoint ordering approach, designed to handle object symmetries that otherwise cause inconsistencies in keypoint assignments. Previous keypoint-based methods relied on fixed keypoint orderings, which failed to account for the multiple valid configurations exhibited by symmetric objects, which our ordering approach exploits to enhance the model’s ability to learn stable keypoint representations. Extensive experiments on the benchmark LINEMOD, Occlusion LINEMOD and YCB-Video datasets show that DLTPose outperforms existing methods, especially for symmetric and occluded objects, demonstrating superior Mean Average Recall values of 86.5% (LM), 79.7% (LM-O) and 89.5% (YCB-V). The code is available at https://anonymous.4open.science/r/DLTPose_/.
@@ 1 Introduction
Among modern approaches, two dominant paradigms have emerged: sparse and dense methods [s24041076]. Sparse methods  [Wu, He, Peng] focus on predicting a small set of keypoints with high accuracy, which ultimately leads to increased accuracy in estimated poses. While they can be highly accurate, one limitation of sparse methods is that their reliance on a limited number of keypoints makes them more susceptible to occlusions, as missing or misidentified keypoints can significantly impact overall estimation. In contrast, dense methods  [Brachmann, Park, Haugaard, labbe2020] predict per-pixel 3D surface coordinates in the object frame using 2D image inputs, providing high robustness through redundant predictions. This redundancy helps mitigate errors by utilizing techniques such as RANSAC to filter out inconsistent and inaccurate predictions. While the increased number of surface points estimated can increase accuracy in occluded and cluttered scenes, dense methods do not specifically focus on obtaining highly accurate estimates of the 3D visible surface of the object. Instead, their primary objective is to generate a broad and redundant representation, often at the expense of accurately localizing individual points.
@@
This work proposes a unified approach that combines the accuracy of sparse with the redundancy of dense methods. Our method, called DLTPose, trains a CNN on RGB-D data to estimate a per-pixel (minimally) four-dimensional representation, where each channel encodes a radial distance, defined as the Euclidean distance between a 3D scene point corresponding to a 2D image pixel, relative to a predefined 3D keypoint. A minimum of four keypoints is required to estimate the pixel surface coordinates in the object-centric reference frame, by solving a novel and highly accurate Direct Linear Transform (DLT) formulation. In this way, the radial quantities inferred from the network are combined with the DLT method to produce an accurate 2D-3D estimate of the visible surface of the object. The estimated object frame points, along with their corresponding 3D camera frame points, are then passed to a RANSAC-enabled Umeyama algorithm [umeyama] to estimate the final object pose. This pipeline enables precise and robust dense 3D surface estimation, ultimately leading to improved pose accuracy compared to previous methods where the dense 2D-3D surface points are inferred using end-to-end networks, albeit with lower accuracy [Park, labbe2020, Haugaard].
@@
By integrating the precision of sparse keypoint-based methods with the redundancy of dense approaches, the proposed framework enhances both robustness and accuracy in object pose estimation. This unified strategy improves the fidelity of 3D surface reconstruction, which in turn leads to more reliable pose estimation, particularly in occluded scenarios where conventional methods struggle.
@@
Handling objects with inherent symmetries is challenging, as visually identical orientations can mislead network training and prediction. Standard loss functions often fail to address these symmetries, resulting in large errors when equivalent poses are treated as distinct. Previous methods attempt to mitigate this by restricting training poses [kehl, Markus] or mapping predictions to the nearest symmetric equivalent [Mahdi], but these approaches struggle with discrete symmetries, leading to reduced generalization and inconsistencies. Instead of limiting training diversity or applying post-hoc corrections, our approach inherently incorporates symmetry awareness by ensuring keypoints remain stable across symmetrical transformations. Unlike traditional keypoint-based methods, which assign fixed keypoints and suffer from inconsistencies under symmetric rotations, our method maintains stable keypoint relationships, improving pose accuracy and robustness across different viewpoints.
@@ 3.1 DLT Surface Estimation
$\begin{bmatrix}\!\!-2\overline{x}_{k_{1}}\!\!\!\!\!\!\!\!&\!\!\!\!-2\overline{y}_{k_{1}}\!\!\!\!\!\!\!\!&\!\!\!\!\!\!\!\!-2\overline{z}_{k_{1}}\!\!\!\!\!\!\!\!&\!\!\!\!1&\!\!\!\!(|\!|\overline{k}_{1}|\!|^{2}\!-\!\widehat{r}_{1}^{2})\\ \vdots&\vdots&\vdots&\!\!\!\!\vdots&\vdots\\ \!-2\overline{x}_{k_{N_{k}}}\!\!\!\!\!\!\!\!&-2\overline{y}_{k_{N_{k}}}\!\!\!\!\!\!\!\!&-2\overline{z}_{k_{N_{k}}}\!\!\!\!&\!\!\!\!1&\!\!\!\!(|\!|\overline{k}_{N_{k}}|\!|^{2}\!-\!\widehat{r}_{N_{k}}^{2})\\ \end{bmatrix}\begin{bmatrix}\overline{x}\\ \overline{y}\\ \overline{z}\\ |\!|\overline{p}|\!|^{2}\\ 1\end{bmatrix}=0.$ (2)
@@ 3.2 Network Architecture
Unlike [wang2019normalized], we discretize the normalized coordinate values to fall within a coarse discrete range $[0,\ldots,n_{\text{b}}-1]$ for each of the three dimensions of normalized space. We found in practise that this discretization served to improve performance for nearly symmetric objects (such as the LINEMOD “glue” object), but not for truly symmetric objects (such as “eggbox”) which suffered from the regular ambiguities, as demonstrated in the ablation study in the Supplementary Material (Sec. S.5). For this reason, we call $\mathcal{L}_{P}$ the pseudo-symmetric loss. This loss term ensures that nearly symmetric objects are correctly aligned under the set $\Theta$ of pseudo-symmetric rotations, which improves the accuracy of radial regression.
@@ 3.3 Symmetric Keypoints
An example is shown in Fig. 3, in which the object exhibits a $\pi$-radian symmetry between pairs of keypoints $k_{1}$$k_{2}$ and $k_{3}$$k_{4}$, yielding a total of 4 possible orderings. For the configuration in Fig. 3(a), the four-channel radial map ($\mathbf{R}_{S}$ of Fig. 2) is ordered as $[\mathbf{R}_{S}^{2},\mathbf{R}_{S}^{1},\mathbf{R}_{S}^{4},\mathbf{R}_{S}^{3}]$ with the order determined by the keypoints’ relative proximity to the camera, i.e. $k_{2}$ is closer to the camera origin than $k_{1}$, and $k_{4}$ is closer than $k_{3}$. For Fig. 3(b), the correct ordering is $[\mathbf{R}_{S}^{1},\mathbf{R}_{S}^{2},\mathbf{R}_{S}^{3},\mathbf{R}_{S}^{4}]$ whereas in Fig. 3(d), the order is $[\mathbf{R}_{S}^{1},\mathbf{R}_{S}^{2},\mathbf{R}_{S}^{4},\mathbf{R}_{S}^{3}]$ while Fig. 3(e) follows order $[\mathbf{R}_{S}^{2},\mathbf{R}_{S}^{1},\mathbf{R}_{S}^{3},\mathbf{R}_{S}^{4}]$. Ensuring a consistent radial map order relative to the camera viewpoint prevents inconsistencies in loss computation, stabilizing training and improving pose estimation under symmetric transformations, as demonstrated in the ablation experiments of Sec. 4.5.
@@ 4.1 Datasets and Evaluation Metrics
Our primary evaluation metric is Average Recall ($AR$[hodan2018bop], which evaluates pose estimation performance across three key components: Visible Surface Discrepancy ($AR_{VSD}$), Maximum Symmetry-Aware Surface Distance ($AR_{MSSD}$), and Maximum Symmetry-Aware Projection Distance ($AR_{MSPD}$). These components enable a fine-grained analysis of pose accuracy while effectively handling object symmetries. For completeness, we also report results using the widely adopted ADD(-S) [hinterstoisser2012model] metric for the LM and LM-O datasets, as well as the ADD-S and ADD(-S) AUC [xiang2018posecnn] metric for the YCB-V dataset.
Method $\textit{AR}_{VSD}$ $\textit{AR}_{MSSD}$ $\textit{AR}_{MSPD}$ Mean AR
LM LM-O YCB-V LM LM-O YCB-V LM LM-O YCB-V LM LM-O YCB-V
PVNet [Peng] 0.502 0.683 0.730 0.638
ZebraPose [su2022zebrapose] 0.598 0.831 0.800 0.903 0.859 0.864 0.752 0.866
CosyPose [labbe2020] 0.670 0.567 0.831 0.810 0.748 0.903 0.849 0.826 0.850 0.773 0.714 0.861
RCVPose [Wu] 0.740 0.682 0.857 0.826 0.773 0.861 0.832 0.792 0.859 0.799 0.749 0.859
Pix2Pose [Park] 0.473 0.766 0.631 0.817 0.659 0.758 0.588 0.780
SurfEmb [Haugaard] 0.615 0.757 0.809 0.849 0.856 0.792 0.760 0.799
PFA [hu2022pfa] 0.658 0.863 0.843 0.920 0.890 0.881 0.797 0.888
CIR [lipson2022coupled] 0.601 0.871 0.778 0.924 0.824 0.885 0.734 0.893
DLTPose (Ours w/o ICP) 0.766 0.578 0.723 0.848 0.731 0.835 0.862 0.791 0.728 0.825 0.700 0.762
DLTPose (Ours) 0.801 0.678 0.879 0.891 0.834 0.922 0.903 0.877 0.884 0.865 0.797 0.895
Table 1: Performance comparison of leading 6DoF PE methods using Average Recall metrics on the LM, LM-O, and YCB-V datasets.
@@ 4.3 Results
Table 1 compares DLTPose with recent state-of-the-art methods on LM, LM-O, and YCB-V using Average Recall ($AR$). All methods refine initial estimated poses with a post processing step, typically ICP. Our method achieves the highest mean $AR$ across all datasets, with scores of 0.865 on LM, 0.797 on LM-O, and 0.895 on YCB-V.
For LM, DLTPose achieves 0.801 in $AR_{VSD}$, surpassing RCVPose (0.740) by +8.2% and CosyPose (0.670) by +19.5%. Additionally, our method obtains 0.891 in $AR_{MSSD}$, outperforming RCVPose (0.826) by +7.8% and CosyPose (0.810) by +10.0%. DLTPose also leads in $AR_{MSPD}$ with 0.903, improving over CosyPose (0.849) by +6.4% and RCVPose (0.832) by +8.5%. Consequently, our mean $AR$ score (0.865) outperforms RCVPose (0.799) by +8.3% and CosyPose (0.773) by +11.9%, establishing DLTPose as the current top-performing method on LM.
For LM-O, DLTPose achieves the highest Mean AR (0.797), outperforming PFA (0.797) and CIR (0.734) by +8.6%, demonstrating superior overall stability in occlusion-heavy settings. It also ranks second across $AR_{VSD}$ (0.678), $AR_{MSSD}$ (0.834), and $AR_{MSPD}$ (0.877). For YCB-V, DLTPose achieves the highest overall mean AR (0.895), surpassing CIR (0.893) by +0.2% and PFA (0.888) by +0.8%. While CIR leads in $AR_{MSPD}$ (0.885), slightly ahead of DLTPose (0.884) by -0.1%, our method ranks first in $AR_{VSD}$ (0.879), outperforming RCVPose (0.857) by +2.6%, and first in $AR_{MSSD}$ (0.922), surpassing CIR (0.924) by -0.2%.
Method LM LM-O YCB-V
-S (-S)
DenseFusion [wang2019densefusion] 94.3
REDE [Hua2020REDEEO] 98.9 65.4
MaskedFusion [pereira2020maskedfusion] 97.3
PR-GCN [Zhou2021PRGCNAD] 99.6 64.8
EANet [ea-net] 97.6 94.2
PVN3D [He] 99.4 63.3 96.1 92.3
RCVPose [Wu] 99.7 71.1 97.2
IRPE [Jin2024IRPEIR] 98.8 85.4 94.9 90.6
PoseCNN [xiang2018posecnn] 78.0 93.0 79.3
ZebraPose [su2022zebrapose] 78.5 92.0 87.5
DLTPose (Ours, w/o ICP) 99.0 80.3 98.3 93.9
DLTPose (Ours) 99.9 90.4 99.7 97.1
Table 2: Performance comparison of leading 6DoF PE methods using the ADD-S and ADD(-S) metrics, on the LM, LM-O, and YCB-V datasets.
In Table 2, we provide the mean ADD(-S) for LM and LM-O across all objects in the dataset for each approach, as well as the AUC for ADD-S and ADD(-S) on YCB-V. Detailed per-object results are provided in the Supplementary Material (Tables  S.3 – S.5). These results show that DLTPose consistently delivers top-tier performance across key metrics while maintaining stability across object categories.
4.4 Surface Estimation Accuracy Experiments
@@
We conducted two experiments that show the improved accuracies and densities of surface points estimated with our DLT method, and its benefit on pose estimation. The first experiment compares the accuracies and number of surface points generated by leading dense approach SurfEmb [Haugaard] with our DLT approach, comparing them against ground truth values. Ground truth 2D-3D correspondences are obtained from the 3D mesh model, foreground mask, and object pose for each object in the LM-O dataset. From the known object pose, the 3D mesh is transformed into the camera frame. Each foreground point in the transformed mesh is then projected onto the image plane using the camera intrinsics, producing a ground truth set in which each 2D image pixel indexes a corresponding 3D object surface point.
@@
Figure 4: Mean error vs. mean number of estimated surface points per object, for SurfEmb (red) and DLTPose (green). Solid lines average over all estimated points; dashed lines only include points with errors $\leq 10$ mm.
@@
Figure 5: Average recall components (MSSD, MSPD, and $AR$) vs. standard deviation of injected noise on the estimated LM-O object surface points, both with (solid lines) and without (dotted lines) ICP refinement.
A second experiment showed the impact of higher accuracy surface estimates on pose estimation. Zero-mean Gaussian noise with varying standard deviations was added to the estimated object surface points, and the $AR$ metric was evaluated for all objects in the LM-O dataset following pose estimation, with the frontend RANSAC-enabled Umeyama used during inference. The results are plotted in Figure 5, and show that as noise increases, performance degrades across evaluation metrics MSSD, MSPD, and $AR$, reinforcing the impact of accurate surface point estimation on the accuracy of pose estimation.

### v2

@@ Abstract
We propose DLTPose, a novel method for 6DoF object pose estimation from RGBD images that combines the accuracy of sparse keypoint methods with the robustness of dense pixel-wise predictions. DLTPose predicts per-pixel radial distances to a set of minimally four keypoints, which are then fed into our novel Direct Linear Transform (DLT) formulation to produce accurate 3D object frame surface estimates, leading to better 6DoF pose estimation. Additionally, we introduce a novel symmetry-aware keypoint ordering approach, designed to handle object symmetries that otherwise cause inconsistencies in keypoint assignments. Previous keypoint-based methods relied on fixed keypoint orderings, which failed to account for the multiple valid configurations exhibited by symmetric objects, which our ordering approach exploits to enhance the model’s ability to learn stable keypoint representations. Extensive experiments on the benchmark LINEMOD, Occlusion LINEMOD and YCB-Video datasets show that DLTPose outperforms existing methods, especially for symmetric and occluded objects. The code is available at https://anonymous.4open.science/r/DLTPose_/.
@@ 1 Introduction
Among modern approaches, two dominant paradigms have emerged: sparse and dense methods [s24041076]. Sparse methods  [Wu, He, Peng] focus on predicting a small set of keypoints with high accuracy, which ultimately leads to increased accuracy in estimated poses. While they can be highly accurate, one limitation of sparse methods is that their reliance on a limited number of keypoints makes them more susceptible to occlusions, as missing or misidentified keypoints can significantly impact overall estimation. In contrast, dense methods  [Brachmann, Park, Haugaard, labbe2020] predict per-pixel 3D surface coordinates in the object frame using 2D image inputs, providing high robustness through redundant predictions. This redundancy helps mitigate errors by utilizing techniques such as RANSAC to filter out inconsistent and inaccurate predictions. While the increased number of surface points estimated can increase accuracy in occluded and cluttered scenes, dense methods have not specifically focused on obtaining highly accurate estimates of the 3D visible surface of the object. Instead, their primary objective has been to generate a broad and redundant representation, often at the expense of accurately localizing individual points.
@@
This work proposes a unified approach that combines the accuracy of sparse with the redundancy of dense methods. Our method, called DLTPose, trains a CNN on RGB-D data to estimate a per-pixel (minimally) four-dimensional representation, where each channel encodes a radial distance, defined as the Euclidean distance between a 3D scene point corresponding to a 2D image pixel, relative to a predefined 3D keypoint. A minimum of four keypoints is required to estimate the pixel surface coordinates in the object-centric reference frame, by solving a novel and highly accurate Direct Linear Transform (DLT) formulation. In this way, the radial quantities inferred from the network are combined with the DLT method to produce an accurate 2D-3D estimate of the visible surface of the object. The estimated object frame points, along with their corresponding 3D camera frame points, are then passed to a RANSAC-enabled Umeyama algorithm [umeyama] to estimate the final object pose.
@@
This pipeline enables precise and robust dense 3D surface estimation, leading to improved pose accuracy compared to previous methods where dense 2D-3D surface points are inferred using end-to-end networks, albeit with lower accuracy [Park, labbe2020, Haugaard]. By integrating the precision of sparse keypoint-based methods with the redundancy of dense approaches, the proposed framework enhances both robustness and accuracy in object pose estimation. This unified strategy improves the fidelity of 3D surface reconstruction, which in turn leads to more reliable pose estimation, particularly in occluded scenarios where conventional methods struggle.
@@
Handling objects with inherent symmetries is challenging, as visually identical orientations can mislead network training and prediction. Standard loss functions often fail to address these symmetries, resulting in large errors when equivalent poses are treated as distinct. Previous methods attempt to mitigate this by restricting training poses [kehl, Markus] or mapping predictions to the nearest symmetric equivalent [rad2017bb8], but these approaches struggle with discrete symmetries, leading to reduced generalization and inconsistencies. Rather than limiting training diversity or applying post-hoc corrections, our approach inherently incorporates symmetry awareness by ensuring keypoints remain stable across symmetrical transformations. Unlike traditional keypoint-based methods, which assign fixed keypoints and suffer from inconsistencies under symmetric rotations, our method maintains stable keypoint relationships, improving pose accuracy and robustness across different viewpoints.
@@ 3.1 DLT Surface Estimation
[…]
@@ 3.2 Network Architecture
Unlike [wang2019normalized], we discretize the normalized coordinate values to fall within a coarse discrete range $0,\ldots,n_{\text{b}}-1]$ for each of the three dimensions of normalized space. We found in practise that this discretization served to improve performance for nearly symmetric objects (such as the LINEMOD “glue” object), but not for truly symmetric objects (such as “eggbox”) which suffered from the regular ambiguities, as demonstrated in the ablation study in the Supplementary Material (Sec. S.5). For this reason, we call $\mathcal{L}_{P}$ the pseudo-symmetric loss. This loss term ensures that nearly symmetric objects are correctly aligned under the set $\Theta$ of pseudo-symmetric rotations, which improves the accuracy of radial regression.
@@ 3.3 Symmetric Keypoints
An example is shown in Fig. 3, in which the object exhibits a $\pi$-radian symmetry between pairs of keypoints $k_{1}$$k_{2}$ and $k_{3}$$k_{4}$, yielding a total of 4 possible orderings. For the configuration in Fig. 3(a), the four-channel radial map ($\mathbf{R}_{S}$ of Fig. 2) is ordered as $\mathbf{R}_{S}^{2},\mathbf{R}_{S}^{1},\mathbf{R}_{S}^{4},\mathbf{R}_{S}^{3}]$ with the order determined by the keypoints’ relative proximity to the camera, i.e. $k_{2}$ is closer to the camera origin than $k_{1}$, and $k_{4}$ is closer than $k_{3}$. For Fig. 3(b), the correct ordering is $\mathbf{R}_{S}^{1},\mathbf{R}_{S}^{2},\mathbf{R}_{S}^{3},\mathbf{R}_{S}^{4}]$ whereas in Fig. 3(d), the order is $\mathbf{R}_{S}^{1},\mathbf{R}_{S}^{2},\mathbf{R}_{S}^{4},\mathbf{R}_{S}^{3}]$ while Fig. 3(e) follows order $\mathbf{R}_{S}^{2},\mathbf{R}_{S}^{1},\mathbf{R}_{S}^{3},\mathbf{R}_{S}^{4}]$. Ensuring a consistent radial map order relative to the camera viewpoint prevents inconsistencies in loss computation, stabilizing training and improving pose estimation under symmetric transformations, as demonstrated in the ablation experiments of Sec. 4.4.
@@ 4.1 Datasets and Evaluation Metrics
Our primary evaluation metric is Average Recall (ARAR[hodan2018bop], which evaluates pose estimation performance across three key components: Visible Surface Discrepancy ($AR_{VSD}$), Maximum Symmetry-Aware Surface Distance ($AR_{MSSD}$), and Maximum Symmetry-Aware Projection Distance ($AR_{MSPD}$). These components enable a fine-grained analysis of pose accuracy while effectively handling object symmetries. For completeness, we also report results using the widely adopted ADD(-S) [hinterstoisser2012model] metric for the LM and LM-O datasets, as well as the ADD-S and ADD(-S) AUC [xiang2018posecnn] metric for the YCB-V dataset.
Method Depth Refinement 1 Model Data ADD(-S) AUC(-S) Mean AR
per Object syn real LM LM-O YCB-V LM LM-O YCB-V
PoseCNN [xiang2018posecnn] 78.0 79.3
PVNet [Peng] 86.3 40.8 73.4 0.575
PVN3D [He] 99.4 63.3 92.3
ZebraPose [su2022zebrapose] 76.9 85.3 0.718 0.815
CosyPose [labbe2020] 84.5 0.633 0.821
RCVPose [Wu] 99.7 71.1 97.2 0.799 0.749 0.859
Pix2Pose [Park] 72.4 32.0
FFB6D [he2021ffb6d] 99.7 66.2 92.7 0.773 0.687 0.758
IRPE [Jin2024IRPEIR] 98.8 85.4 90.6
SurfEmb [Haugaard] 0.758 0.824
HiPose [lin2024hipose] 89.6 95.5 0.799 0.907
EPRO-GDR [10896093] 88.7 92.6 0.786 0.844
CIR [lipson2022coupled] 0.734 0.893
DFTr [zhou2023deep] 99.8 77.7 94.8
DLTPose (Ours, w/o ICP) 99.0 83.6 94.3 0.825 0.714 0.762
DLTPose (Ours) 99.9 91.0 97.4 0.865 0.800 0.909
Table 1: Comprehensive comparison of DLTPose with recent state-of-the-art methods on the LM, LM-O, and YCB-V datasets. Along with accuracy metrics, we report the use of depth input, post-refinement, per-object model training, and the type of training data.
@@ 4.3 Results
Table 1 presents a comprehensive comparison of DLTPose with recent state-of-the-art methods on the LM, LM-O, and YCB-V datasets across multiple evaluation metrics. In addition to performance metrics, the table also reports whether or not each method uses depth input, applies refinement, trains separate models per object, and the type of data used during training. DLTPose achieves the highest mean performance across all three datasets, with a mean AR of 0.865 on LM, 0.800 on LM-O, and 0.909 on YCB-V, while also leading in ADD(-S) and AUC(-S) metrics. Detailed per-object results are provided in the Supplementary Material (Tables S.9S.11).
Methods being compared have pushed the boundaries of 6DoF object pose estimation using RGB-D inputs, yet they exhibit key limitations that DLTPose addresses. DFTr [zhou2023deep] and IRPE [Jin2024IRPEIR] utilize transformer-based fusion or geometric reasoning, achieving high accuracy but relying heavily on vector-wise voting and multi-stage pipelines, which increase training and inference complexity. HiPose’s [lin2024hipose] discrete binary surface codes and iterative pruning introduce non-differentiable control flow, scattering errors across stages and obscuring gradient paths, which makes training behavior difficult to trace and debug. Probabilistic methods like EPRO-GDR [10896093] output uncertainty-aware pose distributions, but sampling introduces gradient noise that destabilizes training and obscures error sources. SurfEmb [Haugaard], which uses dense surface embeddings similar to our method, nevertheless requires refinement and struggles with symmetric ambiguities. In contrast, DLTPose offers a straightforward solution that regresses dense radial distances to canonical keypoints, leverages a novel DLT-based pose solver, and incorporates a symmetry-aware keypoint ordering mechanism to achieve superior accuracy, simplicity, and robustness.
Runtime Analysis. Inference experiments were conducted on a system with an Intel Core i9 13th Gen CPU and a single NVIDIA P100 GPU. Our average object pose estimation time across the LM, LM-O, and YCB-V datasets is 0.627 seconds per image. The average 2D detection time with Mask R-CNN [maskrcnn] on those three datasets is 0.067 seconds.
@@
Surface Estimation Accuracy Experiments. We conducted two experiments that show the improved accuracies and densities of surface points estimated with our DLT method, and its benefit on pose estimation. The first compares the accuracies and number of surface points generated by leading dense approach SurfEmb [Haugaard] with our DLT approach, comparing them against ground truth values. Ground truth 2D-3D correspondences are obtained from the 3D mesh model, foreground mask, and object pose for each object in the LM-O dataset. From the known object pose, the 3D mesh is transformed into the camera frame. Each foreground point in the transformed mesh is then projected onto the image plane using the camera intrinsics, producing a ground truth set in which each 2D image pixel indexes a corresponding 3D object surface point.
@@
Figure 4: Mean error vs. mean number of estimated surface points per object, for SurfEmb (red) and DLTPose (green). Solid lines average over all estimated points; dashed lines only include points with errors $\leq 10$ $m$.
@@
A second experiment showed the impact of higher accuracy surface estimates on pose estimation. Zero-mean Gaussian noise with varying standard deviations was added to the estimated object surface points, and the $AR$ metric was evaluated for all objects in the LM-O dataset following pose estimation, with the frontend RANSAC-enabled Umeyama used during inference. The results are plotted in Figure S.1, and show that as noise increases, performance degrades across evaluation metrics MSSD, MSPD, and $AR$, reinforcing the impact of accurate surface point estimation on the accuracy of pose estimation.